OpenStudy (anonymous):

how many license plates are possible if two letters are followed by four digits?

Parth (parthkohli):

That's the same type of question, sire.

OpenStudy (anonymous):

thought parth just did this one

Parth (parthkohli):

\( \color{Black}{\Rightarrow 10 \times 10 \times 26 \times 26 \times 26 \times 26}\)

OpenStudy (anonymous):

okkk

Parth (parthkohli):

Hmm. The last one had four letters and then 2 digits.

Parth (parthkohli):

I was about to report spam right there.

OpenStudy (anonymous):

so 26x26x10x10x10x10

Parth (parthkohli):

Oh yeah. I got confused up there

OpenStudy (anonymous):

k

Parth (parthkohli):

That magical letter that blows up everyone's mind ^